2. You roll two fair, six-sided dice. Let X be the number on the first die. Let Y be the number on the second die. CalculateThe final answer is 4.472
Answer #1

\small \text{Since the dices are fair and outcome of two dices are independent of each other, the joint distribution} \\ \text{of } X \text{ and } Y \text{ is given as},\\ P(X = i, Y = j) = P(X = i)*(Y = j) = \frac{1}{6}*\frac{1}{6} = \frac{1}{36}; \ \ \ i,j \in \{1,2,3,4,5,6 \}\small \text{Using LOTUS, we get then} \\ \begin{align*} E[\max(X, Y)] &= \sum_{i=1}^6 \sum_{j=1}^6 \max(i,j)*P(X = i, Y = j)\\ &= \sum_{i=1}^6 \sum_{j=1}^6 \max(i,j)*\frac{1}{36} \\ &= \frac{1}{36} \sum_{i=1}^6 \sum_{j=1}^6 \max(i,j) \\ &= \frac{1}{36} \left( \sum_{j=1}^6 \max(1, j) + \sum_{j=1}^6 \max(2, j) + \sum_{j=1}^6 \max(3, j) + \sum_{j=1}^6 \max(4, j) + \sum_{j=1}^6 \max(5, j) + \sum_{j=1}^6 \max(6, j)\right ) \\ &= \frac{1}{36}((1+2+3+4+5+6) + (2+2+3+4+5+6) + (3+3+3+4+5+6) + \\ & \ \ \ \ \ \ \ \ \ \ (4+4+4+4+5+6) + (5+5+5+5+5+6) + (6+6+6+6+6+6)) \\ &= \frac{161}{36} \\ &= 4.472 \end{align*}